OpenStudy (anonymous):

Solve: –(3x – 2) = –4(x + 1) – 4.....options= -10, 10, -2, 2.

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

-3x+2=-4x-4-4 -3x+4x=-8-2 x=-10

OpenStudy (anonymous):

ahh